Is (x-3) a factor of p(x)=x^4-5x^2-27x+15?

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

No.
If you sub 3 for x, p(3) <> 0

Answer by ewatrrr(24785)

 
Hi
The factor theorem states that a polynomial f(x) has a factor (x − k) if and only if f(k) = 0
p(3)=x^4-5x^2-27x+15 = 81-45-81+15 = -30 (x-3) is NOT a factor